0.2x + 0.4y = 1.7
8.3x - 6.3y = -4.3

Answer by Alan3354(69443)

0.2x + 0.4y = 1.7
8.3x - 6.3y = -4.3
-----------
x + 2y = 8.5
83x - 63y = -43
==================
83x + 166y = 705.5
83x - 63y = -43
---------------- Subtract
229y = 748.5
y = 1497/458
-------------
x = 8.5 - 2y
x = 899/458